# Enumerate–Conjecture–Prove: Formally Solving Answer-Construction Problems in Math Competitions

### v3

@@ Abstract
Mathematical reasoning lies at the heart of artificial intelligence, underpinning applications in education, program verification, and research-level mathematical discovery. Mathematical competitions, in particular, present two challenging problem types: theorem-proving, which requires rigorous proofs of stated conclusions, and answer-construction, which involves hypothesizing and formally verifying mathematical objects. Large Language Models (LLMs) effectively generate creative candidate answers but struggle with formal verification, while symbolic provers ensure rigor but cannot efficiently handle creative conjecture generation. We introduce the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic method integrating LLM-based enumeration and pattern-driven conjecturing with formal theorem proving. We present ConstructiveBench, a dataset of 3,431 answer-construction problems in various math competitions with verified Lean formalizations. On the ConstructiveBench dataset, ECP improves the accuracy of answer construction from a Chain-of-Thought (CoT) baseline of 14.54% to 45.06% with the gpt-4.1-mini model. Moreover, combined with ECP’s constructed answers, the state-of-the-art DeepSeek-Prover-V2-7B model generates correct proofs for 858 of the 3,431 constructive problems in Lean, achieving 25.01% accuracy, compared to 9.86% for symbolic-only baselines. Our code and dataset are publicly available at https://github.com/JackSun200312/ECP.
1University of Toronto  2Vector Institute  3Georgia Institute of Technology  4Google DeepMind
@@ 1 Introduction
Mathematical reasoning is fundamental to artificial intelligence (Newell and Simon, 1956), enabling significant progress in domains such as mathematics education, formally verified software, and mathematical research itself (Li et al., 2024). Recent advancements show that LLMs, when guided by careful prompting and tool integration, excel in solving straightforward high-school mathematics problems, but substantial challenges remain in problems requiring formal proof or the construction and verification of complex mathematical objects (Yang et al., 2024). MathArena (Balunović et al., 2025) explicitly highlights this gap, noting that state-of-the-art LLMs achieve near-perfect accuracy on answer-only contests like the American Invitational Mathematics Examination (AIME), yet consistently fail on the United States of America Mathematical Olympiad (USAMO), which demands rigorous proofs and creative answer constructions.
Mathematical competitions, in particular, present two challenging problem types: theorem-proving, requiring rigorous proofs of stated conclusions, and answer-construction, involving hypothesizing and formally verifying mathematical objects. In theorem-proving problems, a clear conclusion is stated, and the goal is a formal proof based on provided hypotheses. Conversely, answer-construction problems require generating specific mathematical objects (e.g., functions, numbers, sets) that satisfy given conditions, along with proofs of correctness. Representative examples of these problems appear in Table 1. Crucially, answer-construction problems pose the unique challenge of first identifying the correct candidate object or structure before verification.
@@
Table 1: Examples of competition-level problems. The left shows an answer-construction problem requiring a candidate answer and proof. The right is a standard theorem-proving problem.
We distinguish between informal mathematics, characterized by intuitive reasoning and natural language arguments, and formal mathematics, represented explicitly in machine-verifiable proof assistants such as Lean (Moura and Ullrich, 2021). While LLMs are effective at creative and exploratory reasoning, they frequently generate unverifiable and incorrect assertions due to hallucinations (Huang et al., 2025). Symbolic reasoning tools, such as SMT solvers (e.g., Z3 (De Moura and Bjørner, 2008)), serve as the logical backbone of automated reasoning tools but cannot feasibly discover closed-form answers in the vast space of candidates.
Inspired by mathematician George Pólya’s problem-solving methodology (Polya, 2014), which emphasizes systematically exploring simpler cases, identifying patterns, and rigorously proving general cases, we introduce the Enumerate–Conjecture–Prove (ECP) framework. Our framework combines the exploratory strengths of LLMs, which enumerate candidate answers through programmatic execution and generalize patterns to form conjectures, with the rigor of formal theorem-proving methods that verify these conjectures. ECP thus effectively addresses the critical limitation of prior approaches by formally solving answer-construction problems, filling a significant gap in the mathematical reasoning landscape.
Our contributions are threefold: (i) a feedback-driven autoformalization pipeline for answer-construction problems in math competitions; (ii) ECP, a modular framework unifying LLM-driven code generation for enumeration, pattern-based conjecture for answer construction, and formal theorem proving; (iii) ConstructiveBench, a curated dataset of 3,431 competition-grade answer-construction problems with verified Lean formalizations and validated through experiments with LLMs and state-of-the-art theorem-proving techniques.
@@ 2 Related Work
Recent neuro-symbolic systems integrate LLMs with formal proof environments. Goedel-Prover (Lin et al., 2025), the DeepSeek-Prover series (Xin et al., 2024; Ren et al., 2025), and InternLM-Math (Ying et al., 2024b) leverage large-scale synthetic formal datasets and iterative training to generate Lean proofs at scale, achieving state-of-the-art results on benchmarks like miniF2F and PutnamBench. Domain-specific systems like AlphaGeometry (Trinh et al., 2024) and LIPS (Li et al., 2025) combine LLM-based intuition with symbolic reasoning to prove Olympiad geometry and inequality theorems, respectively.
@@
Recent work has explored several directions for enhancing the mathematical reasoning capabilities of LLMs. Prompting-based approaches (Wei et al., 2023; Zhang et al., 2025) use carefully crafted prompts to encourage intermediate steps or increase reasoning complexity. Fine-tuning strategies such as MAmmoTH (Yue et al., 2023) and MetaMath (Yu et al., 2024) improve performance by instruction-tuning on math datasets rich in intermediate rationales or diverse reformulations of mathematical problems. Reinforcement learning-based approaches (Luo et al., 2025; Wang et al., 2023; Shao et al., 2024) improve reasoning by aligning step-by-step problem solving with reward signals. Tool-augmented frameworks like PAL (Gao et al., 2023) and Tora (Gou et al., 2023) offer an orthogonal enhancement by generating code as intermediate reasoning steps and outsourcing execution to Python interpreters or external computational tools. SATLM (Ye et al., 2023) extends this framework by using a Z3 symbolic solver to validate logic-based task specifications.
@@
Benchmarks for evaluating LLM mathematical reasoning vary significantly in difficulty and format. Datasets like GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and MathQA (Amini et al., 2019) focus on grade-school and high-school level math with final-answer supervision. In contrast, OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024) include harder problems but still rely on string-matching for evaluation. Formal datasets such as miniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), PutnamBench (Tsoukalas et al., 2024), and ProofNet (Azerbayev et al., 2023) allow machine-verifiable evaluation in Lean or HOL-Light. LeanWorkBook (Ying et al., 2024a) provides large-scale autoformalization from natural language, while Herald (Gao et al., 2024b) auto-informalizes the Mathlib4 corpus back to natural language.
@@ 3.1 Problem Formulation
Problems in math competitions can generally be classified into two categories: theorem-proving tasks and answer-construction tasks. In a theorem-proving task, the problem explicitly presents a set of hypotheses $P$ and a target conclusion $Q$, and the objective is to formally prove that $P$ implies $Q$. Most prior benchmarks and research on AI for formal mathematics, such as MiniF2F (Zheng et al., 2022) and FIMO (Liu et al., 2023), focus on this task format, which aligns with traditional first-order theorem proving.
In contrast to theorem-proving tasks, an answer-construction task asks for constructing mathematical objects that satisfy specified constraints, without an explicitly stated conclusion. We define an answer-construction task by the following components:
@@
We demonstrate the definition with a simple, straightforward example: “Given $a,b\in\mathbb{Z}$ with $b\neq 0$, find all integers $x\in\mathbb{R}$ such that $ax=b$." Here, the context variables are $(a,b)$ with type $\mathbb{Z}\times\mathbb{Z}$ and property $P(a,b):=b\neq 0$. The answer variable is $x$ with type $\mathbb{R}$, the problem predicate is $Q(a,b,x):=ax=b$, and the answer function is $f(a,b)=\{\frac{b}{a}\}$.
In general, the task of answer-construction can be formulated as follows in Lean’s dependent type theory:
@@
[…]
@@
To ensure answers are not only correct but also semantically meaningful, we further require that constructed answers be expressed in canonical form, meaning they must not echo the problem statement’s syntax or structure, and must be in a closed-form, algebraically simplified representation that generalizes across instances. Notably, Lean itself does not enforce canonical representations (e.g., $\mathbb{N}$, $\mathbb{N}$, or more complex yet equivalent expressions are all accepted), and certain LLM-based provers can trivially “cheat” by echoing the statement, e.g., setting $f(a,b)=\{x\mid ax=b\}$ in the above example or answering $\{f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}\mid x(f(x)+f(y))\geq(f(f(x))+y)f(y)\}$ for the example in Table 1. Such trivial substitutions lead to vacuously high proving accuracies if the answer-construction problem is handed straight to the prover. We therefore separate answer construction using tailored LLM prompts to avoid these trivial echoes before formal verification. Finally, the template and requirements for formalizing such answer-construction problems are incorporated into prompts for autoformalization described in Section 3.2 below.
3.2 Autoformalization Pipeline
Autoformalization translates informal, natural language mathematics into machine-verifiable formal statements in proof assistants such as Lean. Due to a scarcity of aligned informal-to-formal mathematical training data, LLMs still frequently fail to produce precise formalizations (Wu et al., 2022a), leading to syntactic and semantic errors.
To address this limitation, we implement a structured autoformalization pipeline that leverages LLMs, formal feedback from the Lean compiler, and targeted retrieval of formal definitions from external documentation sources, as illustrated in Figure 1. The pipeline comprises three core components: (1) autoformalizations with LLMs, (2) Lean knowledge retrieval, and (3) error-guided feedback loops.
Figure 1: Overview of the autoformalization pipeline with error-based feedback and documentation retrieval: An informal statement is first processed by the LLM to produce a formalization. If the result fails either the Lean compilation check or the semantic check by an LLM judge, error messages and relevant Lean documentation are retrieved and fed back to the LLM. The process continues iteratively until the LLM produces a formalization that passes both checks.
Autoformalizations with LLM.
We utilize specialized mathematical LLMs, including deepseek-math-7b-instruct (Shao et al., 2024) and advanced commercial models gpt-4o-mini and deepseek-v3, for autoformalization. To better align these models with formalizations in high-school and undergraduate mathematics, we fine-tune deepseek-math-7b-instruct and gpt-4o-mini on a curated set of 791 paired informal–formal examples drawn from PutnamBench (Tsoukalas et al., 2024) and FIMO (Liu et al., 2023). Since deepseek-v3 does not yet support in-API fine-tuning, we apply a five-shot prompting strategy to guide its output.
Lean Knowledge Retrieval.
LLMs occasionally hallucinate non-existent definitions or misuse current Lean syntax due to their outdated or scarce training data in formalized statements. To mitigate this issue, we construct an external Lean knowledge base comprising 247,826 entries sourced from the official Lean 4.19.0 documentation and the Mathlib (mathlib Community, 2020) repository via LeanDojo (Yang et al., 2023). After focusing on retaining only those definitions most relevant to high-school-level mathematical formalizations (Appendix A), the database is reduced to 44,477 entries. Each entry is represented by embeddings computed with Sentence Transformers (Reimers and Gurevych, 2019) to facilitate the retrieval of relevant definitions using FAISS (Douze et al., 2024). When an LLM encounters a missing or erroneous definition, the system retrieves the top candidate definitions using both FAISS similarity for embedding-based search and edit distance for fuzzy search, which enhances formalization accuracy and reduces hallucination rates.
Error-Guided Feedback Loop.
Generated formalizations undergo immediate compilation checks in Lean, capturing syntactic errors, misuse of definitions, or semantic discrepancies. Each identified error triggers a structured feedback response, guiding the LLM toward correction in subsequent iterations. This iterative loop continues until formalizations compile successfully and satisfy semantic validation criteria evaluated by another LLM judge, which checks semantic equivalence with the original informal statement, proper usage of key definitions, and avoidance of trivial or uninformative answers. Additionally, we employ a sanity checker script to ensure the formalizations adhere to a unified format, ensuring consistency across the formalization process. Once this loop is complete, the subsequent enumerate–conjecture–prove stages take the formal statement as input and begin problem solving.
@@ 3.3 Enumerate–Conjecture–Prove Framework
ECP systematically integrates LLMs’ capabilities for coding and mathematical reasoning with formal theorem proving within the Lean proof assistant. Figure 2 visualizes the workflow of ECP applied to an example problem.
Figure 2: Illustration of the ECP framework applied to a Balkan MO Shortlist problem in ConstructiveBench.
@@
In the initial Enumerate stage, the “enumerator" LLM generates executable Python scripts designed to enumerate explicit examples or special cases of the given mathematical conditions, using the prompts in Appendix B. These enumeration scripts systematically evaluate finite subsets of context variables under computational constraints (timeout of 60 seconds and a maximum of 100 candidate answers) to gather concrete data points and identify initial numerical or structural patterns.
@@
Leveraging insights from the enumerated examples, a “Conjecturer" LLM attempts to solve the problem and construct a closed-form candidate answer in Lean through mathematical reasoning and pattern-finding in enumerated answers. These conjectures propose candidate answers that are both syntactically simple and semantically meaningful, suitable for subsequent rigorous verification. We use a dedicated prompt template in this stage to enforce canonical answer formats and discourage trivial echoes of the original problem statement.
@@
In the final Prove stage, the constructed answer is substituted into the theorem statement, transforming the answer-construction problem into a standard theorem-proving task. The theorem is then verified using specialized LLMs fine-tuned for theorem proving, including DeepSeek-Prover-V2 (Ren et al., 2025), Goedel-Prover (Lin et al., 2025), and Kimina-Prover (Wang et al., 2025). Additionally, we implement baseline symbolic methods as a sequence of automation tactics as follows:
@@
Successful verification by Lean concludes the ECP process.

### v4

@@ Abstract
Mathematical reasoning is central to artificial intelligence, with applications in education, code generation, and research-level mathematical discovery. Mathematical competitions highlight two problem types: theorem-proving, requiring rigorous proofs, and answer-construction, requiring creative generation and formal verification of mathematical objects. Existing research reveals that LLMs can tackle difficult answer-construction tasks but are prone to errors from hallucinations and unverifiable steps, while symbolic methods guarantee rigor but falter in creative answer construction. This raises a key understudied question: how to solve answer-construction problems while preserving both LLM creativity and mathematical rigor? To address this problem, we introduce the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic method integrating LLM-based enumeration and pattern-driven conjecturing with formal theorem proving in Lean, and ConstructiveBench, a dataset of 3,640 formal answer-construction problems from math competitions. ECP is model-agnostic and shows consistent improvements over pure LLM baselines: on the subset of PutnamBench for answer construction, ECP formally solves 6 out of 337 answer-construction problems end-to-end (up from 4 without ECP) using GPT-5 mini and DeepSeek-Prover-V2-7B. On ConstructiveBench, ECP achieves 33.1% end-to-end state-of-the-art accuracy (up from 32.5%), demonstrating its potential to advance formal mathematical reasoning by combining LLM conjecturing with formal verification. Our code and dataset are publicly available at GitHub [^1] https://github.com/sunjia72/ECP and Hugging Face [^2] https://huggingface.co/datasets/sunjia72/ConstructiveBench.
1University of Toronto    2Vector Institute    3Georgia Institute of Technology
@@ 1 Introduction
Mathematical reasoning is fundamental to artificial intelligence (Newell and Simon, 1956), enabling significant progress in domains such as mathematics education, formally verified software, and mathematical research itself (Li et al., 2024). Recent advancements show that LLMs, when guided by careful prompting and tool integration, excel in solving straightforward high-school mathematics problems, but substantial challenges remain in problems requiring formal proof or the construction and verification of complex mathematical objects (Yang et al., 2024). MathArena (Balunović et al., 2025) explicitly highlights this gap, noting that state-of-the-art LLMs achieve near-perfect accuracy on answer-only contests like the American Invitational Mathematics Examination (AIME), yet consistently fail on the United States of America Mathematical Olympiad (USAMO), which demands rigorous proofs and creative answer construction.
Mathematical competitions, in particular, present two challenging problem types: theorem-proving, requiring rigorous proofs of stated conclusions, and answer-construction, involving hypothesizing and formally verifying mathematical objects. In theorem-proving problems, a clear conclusion is stated, and the goal is a formal proof based on provided hypotheses. Conversely, answer-construction problems require the construction of specific mathematical objects (e.g., functions, numbers, sets) that satisfy given conditions, along with proofs of correctness. Representative examples of these problems appear in Table 1. Crucially, answer-construction problems present the unique challenge of identifying the correct candidate object or structure prior to formal verification.
@@
Table 1: Examples of competition-level problems. The left panel shows an answer-construction problem requiring a candidate answer and proof. The right panel shows a standard theorem-proving problem.
We distinguish between informal mathematics, characterized by intuitive reasoning and natural language arguments, and formal mathematics, represented explicitly in machine-verifiable proof assistants such as Lean (Moura and Ullrich, 2021). Autoformalization is an automated process that translates informal problem statements into formal representations. While LLMs are effective at creative and exploratory reasoning, they frequently generate unverifiable and incorrect assertions due to hallucinations (Huang et al., 2025). Symbolic reasoning tools, such as SMT solvers (e.g., Z3 (De Moura and Bjørner, 2008)), serve as the logical backbone of automated reasoning tools but cannot feasibly discover closed-form answers in the vast space of candidates.
Inspired by Pólya’s problem-solving methodology (Polya, 2014), which emphasizes systematically exploring simpler cases, identifying patterns, and rigorously proving general cases, we introduce the ECP framework. Our framework combines the exploratory strengths of LLMs, which enumerate candidate answers through programmatic execution and generalize patterns to form conjectures, with the rigor of formal theorem-proving methods that verify these conjectures. ECP thus effectively addresses the critical limitation of prior approaches by formally solving answer-construction problems, filling a significant gap in the mathematical reasoning landscape.
Our contributions are twofold: (i) ECP, a modular framework that unifies LLM-driven code generation for enumeration, pattern-based conjecture for answer construction, and formal theorem proving in Lean; (ii) ConstructiveBench, an autoformalized dataset of 3640 competition-level answer-construction problems with verified Lean formalizations, validated through experiments with LLMs and state-of-the-art theorem-proving techniques.
@@ 2 Related Work
Recent neuro-symbolic systems integrate LLMs with formal proof environments. Goedel-Prover (Lin et al., 2025), the DeepSeek-Prover series (Xin et al., 2024; Ren et al., 2025), InternLM-Math (Ying et al., 2024b), SeedProver (Chen et al., 2025), Kimina-Prover (Wang et al., 2025) and AlphaProof (Teams, 2024) scale formal data generation and RL-style training to improve Lean proof success on miniF2F, PutnamBench, and related evaluations. Domain-specific systems, including AlphaGeometry (Trinh et al., 2024; Chervonyi et al., 2025) and PyEuclid (Li et al., 2025b) for Euclidean geometry, LIPS for inequalities (Li et al., 2025a), target particular subfields with symbolic search and automation. Beyond theorem proving,  Liu et al. (2025b) introduce a search-based formal problem-solving framework (FPS/D-FPS) with new benchmarks for process-verified solutions.
@@
Beyond formal proof generation, the broader math-reasoning literature explores prompting, learning, and tool use. Prompting-based approaches (e.g., chain-of-thought) encourage explicit intermediate reasoning (Wei et al., 2023). Reinforcement learning has recently become central: GRPO—popularized in Shao et al. (2024)—and large-scale systems like DAPO (Yu et al., 2025a) improve step-wise reasoning via preference/advantage signals at scale. Complementary to policy learning, tool-integrated multi-turn training frameworks such as ReTool (Feng et al., 2025) explicitly train agents to plan, call tools, and verify intermediate computations. Orthogonal “code-as-tool” frameworks, including PAL (Gao et al., 2023) and Tora (Gou et al., 2023), and logic-solver integration such as SATLM (Ye et al., 2023), demonstrate the benefits of executing verifiable intermediate steps.  Liu et al. (2025b) propose a broader formal problem-solving framework (FPS/D-FPS) with new benchmarks for process-verified solutions.
@@
We group evaluation datasets into informal and formal categories. Early works in mathematical evaluations, such as GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and MathQA (Amini et al., 2019) assess numerical/short-answer problem solving in natural language. Follow-up works include harder answer-only suites like OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024) though they still rely on string matching for answer verification. In parallel, formal benchmarks like miniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), PutnamBench (Tsoukalas et al., 2024), ProofNet (Azerbayev et al., 2023), LeanWorkBook (Ying et al., 2024a), Herald (Gao et al., 2024b), and the recent FormalMATH (Yu et al., 2025b) enable machine-verifiable evaluation in Lean/HOL. ProverBench (Xin et al., 2024) contributes a mixed set of 325 formalized problems, including AIME-style items, expanding coverage toward competition math. MathConstruct (Balunović et al., 2025) targets constructive proof tasks. CombiBench (Liu et al., 2025a) focuses on combinatorics in Lean.
@@ 3.1 Problem Formulation
Problems in math competitions can generally be classified into two categories: theorem-proving tasks and answer-construction tasks. In a theorem-proving task, the problem explicitly presents a set of hypotheses $P$ and a target conclusion $Q$, and the objective is to formally prove that $P$ implies $Q$. Prior benchmarks and research on AI for formal mathematics, such as MiniF2F (Zheng et al., 2022) and FIMO (Liu et al., 2023), focus on this theorem-proving task format, which aligns with traditional first-order theorem proving.
In contrast to theorem-proving tasks, an answer-construction task asks for the construction of mathematical objects satisfying specified constraints, without an explicitly stated conclusion. We define an answer-construction task by the following components in the context of Lean dependent type theory:
@@
We demonstrate the definition with a simple, straightforward example: “Given $a,b\in\mathbb{Z}$ with $b\neq 0$, find all $x\in\mathbb{R}$ such that ax=bax=b." Here, the context variables are $(a,b)$ with type $\mathbb{Z}\times\mathbb{Z}$ and property $P(a,b):=a\neq 0$. The answer variable is $x$ with type $\mathbb{R}$, the problem predicate is $Q(a,b,x):=ax=b$, and the answer function is $f(a,b)=\{\frac{b}{a}\}$.
In general, the task of answer-construction can be formulated in Lean’s dependent type theory as follows:
@@
Intuitively, the task is to construct a function $f$ such that for any context parameter $a$ that satisfies property $P$, $f(a)$ precisely captures all and only the valid answers. Variants of format constraints on answers can be incorporated by refining the problem predicate $Q$. For example, to enforce minimality, it is sufficient to define $Q_{\mathit{min}}(a,b):=P(a,b)\wedge\forall c<b,\;\neg P(a,c)$ or use Lean’s built-in notions such as IsLeast or IsGreatest. When the problem admits a single unique solution, $f$ degenerates to $\alpha\to\beta$, and it becomes sufficient to show
@@
To ensure answers are not only correct but also semantically meaningful, we further require that constructed answers be expressed in closed-form canonical format without echoing the problem statement’s syntax or structure. For example, in the following task: theorem test: (x y : $\mathbb{N}$) (hpos : 0 < x $\land$ 0 < y) : x ^ 3 + y ^ 3 = x ^ 2 + 42 * x * y + y ^ 2 $\leftrightarrow$ (x, y) $\in$ answer
[…]
The formal answer-construction task is expressed as a theorem statement with a placeholder or an existentially quantified answer. The autoformalization process will be discussed in Section 4.3 in detail.
@@ 3.2 Enumerate–Conjecture–Prove Framework
ECP enables LLM to perform multi-turn tool calls with Python and Lean in the ReAct manner (Yao et al., 2023). The framework and structured output are implemented using the APPL prompt programming language (Dong et al., 2024), and we include the prompt template in Appendix B. Figure 1 visualizes the workflow of ECP applied to an example answer-construction problem end-to-end.
Figure 1: Illustration of the ECP framework applied to a formalized Balkan MO Shortlist problem in OmniMath dataset (Gao et al., 2024a).
@@
In the initial Enumerate stage, the LLM is given a formal answer-construction task, and the goal is to enumerate the candidate answers by generating Python programs. The model performs multi-turn interaction to refine the program and produce enumeration: {<reasoning_1> <python_call_1> <python_response_1><reasoning_2>…}. For problems with unassigned context variables or infinitely many answers, the model is encouraged to explore small customized setups and return bounded enumerations, guided by the system prompt with constraints of timeout=60s and a maximum of 100 enumerated answers.
@@
In the Conjecture (i.e. answer-construction) stage, given the answer-construction task and the enumerations as hint, the LLM reasons about the problem, generalizes beyond enumeration and attempts to propose a syntactically valid closed-form answer in Lean, following the workflow {<reasoning_1> <lean_call_1> <lean_response_1><reasoning_2> …}. To prevent potential answer “hacking" described in Section  3.1, we design the system prompt to explicitly disallow such illegal answers that lead to trivial or vacuous versions of the Lean theorem statement.
@@
In the final Prove stage, the candidate answer is substituted into the theorem statement, reducing the answer-construction problem into a standard theorem-proving task. Our framework supports invoking theorem-proving models, such as DeepSeek-Prover-V2 (Ren et al., 2025), Kimina-Prover (Wang et al., 2025), and Goedel-Prover (Lin et al., 2025), to generate proofs of the theorem, and Lean verifies whether the proof closes the goal. Additionally, we implement pure symbolic methods as a sequence of automation tactics as follows, which will be used to verify equivalence between constructed and ground-truth answer in evaluation:
@@
Successful Lean verification of the proof concludes the ECP pipeline.